object thrown the vorcity is 0 &the t=2 &the g=-9..8 what's the fainal vorcity ?
1
Expert's answer
2017-09-18T08:48:07-0400
Assuming no air resistance we can use the projectile motion formulas.
v_0=0 m/s.
t=2 s.
g=-9.8 m/s^2 .
The final velocity is
v=v_0+gt=0-9.8(2)= -19.6 m/s.
The sign means that the positive direction is upwards.
Answer: -19.6 m/s.